# Continual-MEGA: A Large-scale Benchmark for Generalizable Continual Anomaly Detection

### v2

@@ 6 Conclusion
We present Continual-MEGA, a new large-scale benchmark for continual anomaly detection (AD), built by integrating multiple public datasets and curating a novel dataset, ContinualAD, to significantly enhance sample volume and diversity. Comprehensive evaluations on Continual-MEGA reveal that there is still substantial room for improvement in existing AD methods, highlighting both the need for further research in continual AD and the effectiveness of the curated ContinualAD dataset. Also, we introduce a novel baseline method that integrates MoE-style adaptor modules, anomaly feature synthesis, and prompt-based feature tuning with CLIP. Our approach achieves state-of-the-art performance on Continual-MEGA, providing a strong baseline for future work. Limitation: Since our evaluation benchmark still has class imbalance, enhancing the benchmark by improving category balance and expanding data diversity will improve the Continual-MEGA benchmark.
7 Acknowledgment
This work was supported by the Institute of Information &\&& Communications Technology Planning &\&& Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligent Graduate School Program (Chung-Ang University) and RS-2022-II220124, Development of Artificial Intelligence Technology for Self-Improving Competency-Aware Learning Capabilities].
Appendix A Continual-MEGA Benchmark Configuration Details
@@ A.1 ContinualAD Dataset
To form the Continual-MEGA benchmark, we propose the ContinualAD dataset which has significantly larger quantities and volumes compared to previous datasets. The ContinualAD dataset consists of a total of 30 classes, as listed in Table A. Table D describes the number of normal and anomaly samples for each class in the ContinualAD dataset. Figure A illustrates sample images from the ContinualAD dataset, which includes diverse scenes captured in different background settings. Moreover, the ContinualAD dataset includes a wide range of object instances within the same class, enabling the evaluation of robustness to intra-class variation.
Table A: Class list of ContinualAD dataset.
       Class Names
Energy-bar Apple Kleenex Ruler Toothpaste
Sunglasses Capsule Flash-drive Band-aid Cucumber
Toothbrush Soap Dollar Pencil Fork
Toy Multi-pen Chopsticks Watermelon Egg
Spoon Calculator Eraser Mango Candy
Mouse Glasses-case Notebook Food-container Cup
Consequently, as shown in Figure 2, the image variance of the ContinualAD dataset is significantly higher than that of existing datasets. The variance value is computed by first calculating the pixel-wise variance across images within each class, and then averaging these values across all classes in the dataset. This indicates that prior datasets mainly consist of highly similar images within each class, limiting their ability to evaluate model performance under diverse conditions. In contrast, the higher variance in ContinualAD facilitates more realistic and challenging evaluation scenarios.
@@
A.2 Training Setting for Continual-MEGA Benchmark
To simulate a low-resource environment where training samples are limited, we adopt a minimal supervision setting in the proposed Continual-MEGA benchmark. Specifically, only 10 normal and 10 anomalous training images are provided per class during both the base and continual learning stages. Table C shows the number of training and test images used for base classes and continual learning classes in each scenario. For continual learning, the classes are partitioned into three task settings, with each task comprising 5, 15, and 30 classes, respectively, to simulate varying levels of incremental difficulty. Following recent trends in AD toward few-/zero-shot and continual learning, we adopt a limited number of training and adaptation samples to better reflect realistic constraints. However, depending on the target AD application, the training setup of our ContinualAD benchmark can be flexibly reconfigured to suit different deployment scenarios. This will be discussed in the Limitation section in more detail.
For the proposed baseline method, hyperparameter tuning was conducted solely on the base classes of Scenario 1 in a lightweight manner. The resulting hyperparameters were uniformly used across all remaining scenarios to ensure fair and consistent evaluation. To evaluate the model performance in a realistic continual learning setting, we avoided scenario-specific hyperparameter tuning on purpose. This design choice aim to reflect practical constraints in real-world deployments, where care tuning for each newly incoming task is often infeasible cha2024hyperparameters .
A.3 Text Prompting Details
We used generalized text prompts to obtain text features that are not specific to any particular domain or class. Table B shows the generalized text prompts used to obtain general text features. It consists of 10 prompts for both normal and anomaly classes, respectively.
Table B: Description of generalized text prompts.
  Prompt No. Normal Prompts Anomaly Prompts
4 A photo of a normal object A photo of an anomalous object
5 This is not an anomaly This is an example of an abnormal object
9 A photo of a standard object A photo of an abnormal object
Table C: Number of training and test samples in each scenario.
  Scenario Stage Train Test
#Normal #Anomaly #Normal #Anomaly
Scenario 1 Base 850 850 71,274 44,301
Continual 600 600 49,543 28,140
Scenario 2 Base 580 580 59,121 35,972
Continual 600 600 60,267 34,281
Scenario 3 Base 580 580 69,788 37,130
Continual 300 300 35,245 17,597
Appendix B Continual-MEGA Benchmark Evaluation Details
This section discusses the deeper analysis of the meaning of evaluation results of various AD methods on our Continual-MEGA Benchmark presented in Section 5 of the paper. We note that the scenario 3 exclude ContinualAD datasets both for Base and New classes of the benchmark. Figure B compares the zero-shot performance of various methods on the MVTec-AD and VisA datasets under two conditions: (1) trained only on the Base classes, and (2) after continual adaptation as defined by our proposed Continual-MEGA benchmark. Notably, our proposed baseline model demonstrates improved generalization, benefiting from both a stronger set of Base classes and the continual adaptation of additional classes. Scenario 2 includes our proposed ContinualAD dataset, under which most methods exhibit improved zero-shot performance after continual learning. In contrast, when ContinualAD is excluded in Scenario 3, most existing methods suffer a degradation in zero-shot generalization after continual learning. Among them, our proposed method shows the smallest performance drop, indicating stronger robustness to continual adaptation compared to other approaches. To further investigate this observation, we refer to the quantitative results from Scenarios 2 and 3, presented in Table E and Table F. These tables provide detailed evaluation metrics corresponding to continual adaptation and zero-shot evaluation results. For easier display of the results, we visualize the results for representative methods of the tables in Figure 5.
Table D: Number of normal and anomaly samples per class in the ContinualAD dataset.
@@ Appendix B Continual-MEGA Benchmark Evaluation Details
Table E: Experimental results on Scenario 2. To evaluate the zero-shot generalization performance of the methods, we excluded the MVTec-AD and VisA classes from training and used them only for evaluation. /$\cdot$/$\cdot$/$\cdot$ / ⋅ / ⋅ denotes Image-AUROC, Pixel-AP and average value. Zero-shot performance on MVTec-AD and VisA is presented in Figure 5. The notation ‘$X$-$Y$ ($Z$ tasks)’ denotes an evaluation setup where the model is initially trained on $X$ base classes, followed by $Y$ continual learning phases, each including $Z$ new tasks. The best-performing results are highlighted in bold.
From the results presented, we observe that anomaly detection (AD) performance has following tendencies: (1) Compared to Scenario 3, overall AD performance improves in Scenario 2 across most methods, showing the effectiveness of the ContinualAD dataset. (2) Continual adaptation using the ContinualAD dataset enhances zero-shot generalizability, as observed in VisA and MVTec. Excluding ContinualAD leads to a consistent drop in performance among prior methods. (3) Our proposed baseline achieves strong and consistent results across all scenarios, showing notable improvements in Scenario 2 and maintaining competitive generalizability in Scenario 3 after continual adaptation. Based on these results, we can reasonably conjecture that our ContinualAD dataset provides valuable information for detecting anomalies in both unseen and continually introduced categories under more challenging scenarios than prior setups. The proposed baseline exhibits robust and consistent performance, setting a meaningful benchmark for future AD methods.
(a) MVTec-AD.
(b) VisA.
@@
Figure B: Image-level AUROC and pixel-level AP performance on MVTec-AD (top) and VisA (bottom) datasets. Each point represents the performance of a method before (•) and after ($\times$) continual learning. Arrows indicate the performance change from the model trained only on base classes to the model trained via continual learning. The continual learning results are averaged over three settings, where each task consists of 5, 10, and 30 New classes, respectively.
Table F: Experimental results on Scenario 3. To verify the effectiveness of the ContinualAD dataset, we excluded it from the training process. /$\cdot$/$\cdot$/$\cdot$ / ⋅ / ⋅ denotes Image-AUROC, Pixel-AP, and average value. The notation ‘$X$-$Y$ ($Z$ tasks)’ denotes an evaluation setup where the model is initially trained on $X$ base classes, followed by $Y$ continual learning phases, each including $Z$ new tasks. The best-performing results are highlighted in bold.
Appendix C Deeper Discussion of the Limitation and Future Research
@@
Regarding the dataset sample configuration, a primary limitation of the proposed benchmark is class imbalance, as sample sizes vary significantly across datasets. In our continual evaluation setup, models that better fit classes having a smaller number of samples would be beneficial to achieve higher performance. While this setting reflects the class imbalance commonly observed in real-world inspection tasks, balancing sample quantities across classes would improve the reliability of AD performance evaluation in the continual setup.
@@
Regarding the Benchmark training and evaluation configuration, the Continual-MEGA benchmark intentionally adopts limited training and adaptation samples to evaluate the effectiveness of recent few-/zero-shot AD methods under both unseen and continual setups, leveraging a significantly larger evaluation set. While our primary focus is evaluation, we expect higher accuracy with increased training data, particularly for the Base set, making detailed analysis across varying training sizes a key direction for future research.

### v3

@@ 6 Conclusion
We present Continual-MEGA, a new large-scale benchmark for continual anomaly detection (AD), constructed by integrating multiple public datasets and curating a novel dataset, ContinualAD, to significantly enhance sample volume and diversity. Comprehensive evaluations on Continual-MEGA reveal that existing AD methods still have substantial room for improvement, highlighting both the need for further research in continual AD and the effectiveness of the curated ContinualAD dataset. To address these challenges, we introduce a novel baseline method integrating MoE-style adapter modules, anomaly feature synthesis, and prompt tuning within the CLIP framework. Our baseline exhibits robust and consistent performance, highlighting a promising direction: the synergy between lightweight CLIP adaptation and feature synthesis is vital for achieving both learning stability and zero-shot generalization within continual frameworks.
Impact Statement
This research contributes to the field of industrial anomaly detection by establishing a large-scale benchmark for continual and zero-shot learning. A primary motivation of our work is to mitigate the computational cost and carbon footprint associated with retraining deep learning models in dynamic production environments. The proposed datasets and methods focus exclusively on product defects and do not involve personal identifiable information or surveillance capabilities. Consequently, we believe this work presents minimal ethical risks while offering tangible benefits for manufacturing sustainability and reproducibility in machine learning research.
Appendix A Continual-MEGA Benchmark Details
@@ A.1 ContinualAD Dataset
Figure A illustrates sample images from the ContinualAD dataset, which includes diverse scenes captured in different background settings. Table A presents the class names along with the numbers of normal and anomaly samples for each class. The red boxes highlight the anomalous regions. Moreover, the ContinualAD dataset provides multiple instances and varied backgrounds even within the same object class, enabling a more comprehensive evaluation of anomaly detection performance compared to previously released datasets.
@@
Table A: Number of normal and anomaly samples per class in the ContinualAD dataset.
A.2 Scenario Overview and Class Distribution
Table B provides an overview of the three evaluation scenarios in the Continual-MEGA benchmark. Scenario 1 uses all seven datasets in the continual stream. Among them, 85 classes are assigned to the base task, and the remaining 60 classes are introduced as new classes. We consider task sizes of 5, 10, and 30 classes (85–5/10/30). Scenario 2 is designed to evaluate zero-shot performance after continual adaptation. To create a held-out target, MVTec-AD and VisA are removed from the training stream. In this case, the base task consists of 58 classes, and 60 additional classes are used as continual new classes. Zero-shot evaluation is then performed on MVTec-AD and VisA. Scenario 3 examines how the newly collected ContinualAD dataset affects zero-shot performance. Here, ContinualAD is excluded from the adaptation stream, while the other datasets are kept as in Scenario 2. The base task again contains 58 classes, but only 30 classes are used as continual new classes. As in Scenario 2, zero-shot performance is evaluated on MVTec-AD and VisA under this reduced-diversity stream.
Figures DE, and F illustrate the detailed class distributions for Scenario 1, 2 and 3 of the Continual-MEGA Benchmark, respectively. We observe a notable imbalance among classes in all scenarios. This imbalance underscores the challenge of maintaining consistent anomaly detection performance, as smaller classes from previous datasets are more prone to being forgotten during adaptation. In particular, Scenario 3 excludes the ContinualAD dataset, leading to fewer samples and a more constrained class distribution compared to Scenario 2. Such variations in class composition and sample availability across the scenarios make them valuable for comprehensively evaluating the robustness and adaptability of continual anomaly detection methods.
  Scenario (Base-New) Base classes New classes stream Zero-shot evaluation
(datasets) (datasets / #classes)
Scenario 1 (85-5/10/30) All 7 Datasets Same as base / 60 classes
Scenario 2 (58-5/10/30) excluding MVTec-AD, VisA Same as base / 60 classes MVTec-AD, VisA
Scenario 3 (58-5/10/30) excluding MVTec-AD, VisA, ContinualAD Same as base / 30 classes MVTec-AD, VisA
Table B: Overview of the three evaluation scenarios in the Continual-MEGA benchmark. For each scenario, we summarize the datasets used for the base classes, the continual (New) class stream, and any datasets that are held out for zero-shot evaluation.
  Scenario Stage Train Test
#Normal #Anomaly #Normal #Anomaly
Scenario 1 Base 850 850 71,274 44,301
Continual 600 600 49,543 28,140
Scenario 2 Base 580 580 59,121 35,972
Continual 600 600 60,267 34,281
Scenario 3 Base 580 580 69,788 37,130
Continual 300 300 35,245 17,597
Table C: Number of training and test samples in each scenario.
Figure B: Overview architecture of inference process. $B$ and NN denote the adapters corresponding to the base classes and the task-specific classes $\{1,2,3,...,N\}$, respectively. During inference, we use $A^{\text{avg}}$, which is the average of the adapter weights trained on the base classes and the NN task adapters.
A.3 Training Setup for Continual-MEGA Benchmark
To simulate a low-resource environment where training samples are limited, we adopt a minimal supervision setting in the proposed Continual-MEGA benchmark. Specifically, only 10 normal and 10 anomalous training images are provided per class during both the base and continual learning stages. Table C shows the number of training and test images used for base classes and continual learning classes in each scenario. For continual learning, the classes are partitioned into three task settings, with each task comprising 5, 15, and 30 classes, respectively, to simulate varying levels of incremental difficulty. Following recent trends in AD toward few-/zero-shot and continual learning, we adopt a limited number of training and adaptation samples to better reflect realistic constraints. However, depending on the target AD application, the training setup of our Continual-MEGA benchmark can be flexibly reconfigured to suit different deployment scenarios. This will be discussed in the Limitations section in more detail.
Figure 4 and Figure B illustrate an overview of the proposed baseline method, depicting the training and inference processes. During training, the model incorporates Mixture-of-Experts adapters to dynamically specialize representations across tasks, while synthetic anomaly feature generation is employed to enrich limited samples and improve adaptation. Additionally, we leverage generalized text prompts to obtain text features that are not specific to any particular domain or class. Table D lists the prompts used to extract these generalized features, consisting of 10 prompts each for normal and anomaly classes. This design helps the model capture more robust, domain-agnostic representations that improve anomaly detection performance across continual learning tasks.
For the proposed baseline method, hyperparameter tuning was conducted solely on the base classes of Scenario 1 in a lightweight manner. The resulting hyperparameters were uniformly used across all remaining scenarios to ensure fair and consistent evaluation. To evaluate model performance in a realistic continual learning setting, we avoided scenario-specific hyperparameter tuning. This design choice aims to reflect practical constraints in real-world deployments, where careful tuning for each newly incoming task is often infeasible (Cha and Cho, 2024).
  Prompt No. Normal Prompts Anomaly Prompts
4 A photo of a normal object A photo of an anomalous object
5 This is not an anomaly This is an example of an abnormal object
9 A photo of a standard object A photo of an abnormal object
Table D: Description of generalized text prompts.
Image Pixel Image Pixel Image Pixel
Only-normal SimpleNet 58.8 6.3 61.3 4.5 57.5 4.5
GeneralAD 51.5 2.6 52.6 1.8 54.4 2.7
HGAD 59.5 5.0 56.1 3.2 55.5 2.7
ResAD 73.3 15.5 69.1 7.8 70.7 15.3
VLM-based MVFA 81.7 32.6 65.8 10.4 70.7 21.2
VCP-CLIP 73.8 25.4 61.0 23.1 61.9 22.5
MediCLIP 73.9 4.5 78.1 8.5 75.3 5.9
Continual UCAD 55.8 1.6 58.1 4.7 56.0 3.6
IUF 60.5 7.4 57.4 4.4 58.5 4.2
$\text{IUF}^{\ast}$ 68.3 13.5 65.8 9.5 63.6 9.5
Ours 83.1 39.0 82.0 35.7 77.8 36.5
Table E: Experimental results of base classes across scenarios. We note that the base classes used in Scenario 2 and Scenario 3 differ, as ContinualAD is included among the base classes in Scenario 2 but not in Scenario 3. The best-performing results are highlighted in bold.
@@ 10 This object meets standard characteristics An abnormality detected in this object
Figure C: Image-level AUROC and pixel-level AP performance on MVTec-AD (a, b) and VisA (c, d) datasets. Each point represents the performance of a method before (•) and after ($\times$) continual learning. Arrows indicate the performance change from the model trained only on Base classes to the model trained via continual learning. The continual learning results are averaged over three settings, where each task consists of 5, 10, and 30 New classes, respectively. Notably, in Scenario 3, where the proposed the ContinualAD dataset is excluded, most methods experience a noticeable drop in continual zero-shot learning (CZSL) performance, highlighting the importance of incorporating ContinualAD for robust generalization.
A.4 Continual-MEGA Benchmark Evaluation Details
This section provides an in-depth analysis of the evaluation results of various AD methods on our Continual-MEGA Benchmark, complementing the main experiments. We note that Scenario 3 excludes the ContinualAD dataset from both the Base and New classes of the benchmark. Figure C of the paper compares the zero-shot performance of various methods on the MVTec-AD and VisA datasets under two conditions: (1) trained only on the Base classes, and (2) after continual adaptation as defined by our proposed Continual-MEGA benchmark.
From the results presented, we observe that anomaly detection (AD) performance has the following tendencies: (1) Compared to Scenario 3, overall AD performance improves in Scenario 2 across most methods, showing the effectiveness of the ContinualAD dataset. (2) Continual adaptation using the ContinualAD dataset enhances zero-shot generalizability, as observed in MVTec-AD and VisA. Excluding ContinualAD leads to a consistent drop in performance among prior methods. (3) Our proposed baseline achieves strong and consistent results across all scenarios, showing notable improvements in Scenario 2 and maintaining competitive generalizability in Scenario 3 after continual adaptation.
Additionally, performing anomaly detection on the baseline categories in our proposed setup is substantially more challenging than in conventional benchmarks, as shown in Table E. Under this more difficult setting, VLM-based methods demonstrate significantly stronger performance compared to approaches explicitly designed for continual learning. This performance gap stems from the inherently limited detection capability of existing continual anomaly detection methods at their initial stage.
Appendix B Implementation Details
We use CLIP with ViT-L/14 (Dosovitskiy et al., 2020) architecture, which consists of 24 sublayers divided into four layers, where each layer contains six sublayers. The size of input images was set to 336. The adaptation layers for anomaly feature generation were applied to layers 1, 2, 3, and 4. The batch size was set to 16. The adapter parameters are optimized using AdamW (Loshchilov, 2017) with a learning rate of $1\times 10^{-4}$ for 50 epochs on the $Base$ classes and for 20 epochs for each continual task under the same optimization setup. The learnable text-prompt module is trained only on the $Base$ classes using Adam (Kingma and Ba, 2015) with a learning rate of $1\times 10^{-4}$ and is kept frozen during all continual tasks. For synthetic anomaly feature generation, we use the random noise term $\gamma$ as Gaussian noise with standard deviation $0.25$, i.e., $\gamma\sim\mathcal{N}(0,0.25^{2})$, and add it to the feature space.
Appendix C Discussions: Limitations and Future Work
@@
Regarding the dataset sample configuration, a primary limitation of the proposed benchmark is class imbalance, as sample sizes vary significantly across datasets. In our continual evaluation setup, models that better fit classes having a smaller number of samples would be beneficial to achieve higher performance. While this setting reflects the class imbalance observed in real-world inspection, balancing sample quantities across classes would improve the reliability of AD performance evaluation in the continual setup.
@@
Regarding the benchmark training and evaluation configuration, the Continual-MEGA benchmark intentionally adopts limited training and adaptation samples to evaluate the effectiveness of recent few-/zero-shot AD methods under both unseen and continual setups, leveraging a significantly larger evaluation set. While our primary focus is evaluation, we expect higher accuracy with increased training data, particularly for the Base set, making detailed analysis across varying training sizes a key future direction.
@@
Figure D: Class distribution of Scenario 1. This example illustrates Scenario 1 when each task contains 10 classes, and six tasks arrive sequentially. The seven colored background bands indicate one Base block (left) and the six incremental New task blocks (right) that arrive in order. The orange line denotes the anomaly-sample count per class, and the gray bars denote the total sample volume.
Figure E: Class distribution of Scenario 2. This example illustrates Scenario 2 when each task contains 10 classes, and six tasks arrive sequentially. The seven colored background bands indicate one Base block (left) and the six incremental New task blocks (right) that arrive in order. The orange line denotes the anomaly-sample count per class, and the gray bars denote the total sample volume.
Figure F: Class distribution of Scenario 3. This example illustrates Scenario 3 when each task contains 10 classes, and three tasks arrive sequentially. The seven colored background bands indicate one Base block (left) and the six incremental New task blocks (right) that arrive in order. The orange line denotes the anomaly-sample count per class, and the gray bars denote the total sample volume.